OpenStudy (anonymous):

If f(x) = 3x + 5 and g(x) = 2x - 9, find f(x) + g(x)

OpenStudy (anonymous):

f(x)+g(x)=3x+5+2x-9=5x-4

OpenStudy (anonymous):

f(x) + g(x) = (3x + 5) + (2x - 9) = 3x + 5 + 2x - 9 = 5x - 4